Highest comman factor of (21m+4) and (14m+3) is? [where m belongs to N]

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
please post the optio s tooas these question require options too to check. I am solving this as a question which is similiar to this sokved by me in past
 
m = natural number.   Two numbers:  N1: 21 m + 4,     N2:   14 m + 3
1)  N2 is always odd, as 14m is even and 3 is odd. So option a is incorrect. N2 is not divisible by 2. So  2 cannot be HCF.
2)  N1 = 21 m + 4 ,      N2 = 14 m + 3
     Add N1 + N2 =  35 m + 7 = 35 m + 5 + 2
    If N1 and N2 are divisible by 5, then N1+N2 must be divisible by 5.
     But we see clearly that N1+N2 is not divisible by 5.
3)    N1 = 21 m + 4     N2 = 14 m + 3
       N1 - N2 = 7 m + 1 
       3 N2 - 2 N1 = 42 m + 9 - 42 m - 8 = 1
  
    If N1 and N2 are both divisible by 9, then  3 N2 - 2 N1 should be divisible by 9 or it should be 0.  So 9 is not the HCF.